Suppose you travel north for 65 kilometers then travel east 75 kilometers. How far are you from your stating point?
Found 2 solutions by psbhowmick, ptaylor:
Answer by psbhowmick(878)
Answer by ptaylor(2198)

Your route has traced out two legs of a right triangle. The hypotnuse of the right triangle is the shortest distance from where you are to where you started from and this is given by: c^2=a^2+b^2 which is the Pythagorean Theorem. c is the hypotnuse and a and b are the two legs.
c^2=65^2+75^2
c^2=4225+5625
c^2=9850 take square root of both sides
c=99.247 kilometers

Hope this helps----ptaylor